OpenStudy (anonymous):

F(t)=12sint, [0,9π/2]

OpenStudy (anonymous):

what do u want with this?

OpenStudy (anonymous):

?????????

OpenStudy (anonymous):

calculate the avergae value of the function F(t)=12sint, across the interval [0,9π/2] I have the answer but dont understand how to get the end answer

OpenStudy (dumbcow):

\[=\frac{1}{9\pi/2}12\int\limits_{0}^{9\pi/2}\sin t dt\]

OpenStudy (anonymous):

thanks dumb cow :D

OpenStudy (dumbcow):

welcome glad that helped its same formula whenever you need avg value

OpenStudy (anonymous):

sweet :D:D:D will be very useful :D